# An Imitative Reinforcement Learning Framework for Autonomous Dogfight

### v1

@@ 6.1. Comparative Results
According to Table 2, the adaptive weighting method proposed in this study demonstrates superior performance across three key indicators. Specifically, the TD3 method shows the lowest performance in terms of the hit success rate, at $0\%$, among all methods. Conversely, the adaptive weighting method proposed in this study, along with the E-SAC and SAC methods, achieve a hit success rate of $100\%$. However, although the SAC method achieves a success rate of $100\%$, its hit success rate is $0\%$, indicating its failure to learn to correctly shoot down opponent aircraft with missiles. Furthermore, the reward value of the SAC method is significantly lower than that of the E-SAC method, which has mastered the correct method for shooting down enemies. Further comparison between our method and the BC method reveals that, the BC method performs more poorly in terms of both success rate in task completion and reward value compared to our method due to the lack of interaction with the environment and the inability to explore autonomously. It is noteworthy that although BC method has a higher success rate than TD3 method, the average returns of BC method are significantly lower than those of TD3. This is also attributed to the inability to explore autonomously. When compounding errors occur (Xu et al., 2020; Ross and Bagnell, 2010; Ross et al., 2011), BC strategy tends to lose control and thus achieves very low returns, whereas TD3 methods can still closely track the opponent by pursuit strategy, thus preventing a rapid decline in returns.
@@ 6.2. Analytical Results
We show weight $\lambda$ in the two modes (linear and adaptive) in Figure 6, and find that the value of $\lambda$ is related to the learning performance. As shown in Figure 5(a), the linear weighting method proposed in this study demonstrates a faster learning rate in the initial stages of training compared to that of the adaptive weighting method. This is primarily attributed to the higher weight assigned to the imitation of objectives by the linear weighting method in the early stages, accelerating the learning of expert knowledge. However, this approach increases the risk of overfitting and limits the model’s generalization ability, resulting in instability and poor convergence performance in training processes.
@@
Figure 7. The distance between our aircraft and the opponent.
@@
Figure 7 depicts the distance between our aircraft and the opponent in one episode, and the trajectories are collected by the convergent models learned by the corresponding algorithms. Subfigures in Figure 7 annotate the moments of missile launch (red dots), phases of holding missiles (blue background), and stages of successfully locking onto targets (yellow background). If the first red dot on the curve falls within the intersection of the blue and yellow backgrounds, it indicates that the agent has mastered an effective launch policy. Accordingly, it is observed that the method proposed, along with the E-SAC and BC methods, successfully executed missile launches and accurately shot down the opponent aircraft (i.e., mastering the three-dimensional policy of pursuit, lock, and launch). Further comparison of the speed of target lock and missile launch among these three methods reveals that the proposed method achieves both with fewer steps, indicating higher attack efficiency. The distance curves of the TD3 and SAC methods exhibit a downward trend, and there are target locking zones; however, they fail to correctly launch missiles, demonstrating their mastery of only the two-dimensional policy of pursuit and lock. By comparing the data in Table 2, although the SAC method has mastered only a two-dimensional policy, it can successfully shoot down the opponent aircraft through collision. In contrast, the TD3 approach fails to shoot down the opponent aircraft. Therefore, the lock capability of SAC exceeds that of TD3.
Figure 8. The trajectories collected by the learned policy and the opponent.
@@
Figure 8 corresponds respectively to Figure 7, and Figure 8 shows the three-dimensional interaction scenarios between our aircraft and the opponent aircraft using three different methods. The red curves represent the movement trajectories of the opponent aircraft, while the blue curves represent the movement trajectories of our aircraft, with the colors gradually deepening over time. The moments of launch are indicated by red dots on the blue trajectories. Figure 8(a) illustrates the operational effectiveness of the proposed learning framework, i.e., the controlled UCAV swiftly catches up to the opponent aircraft, successfully locking onto it, launching a missile, and accurately hitting the target. By contrast, the UCAV with the TD3 policy (Figure 8(c)) can pursue the opponent aircraft but keeps hovering around it, indicating insufficient lock and launch capabilities. In Figure 7(e), the UCAV employing the SAC policy catches up to, locks onto, and rams the opponent aircraft by collision. Through the above analysis, we systematically summarize the policies mastered by various methods, which are presented in Table 3.
Table 3. Policy analysis
@@
6.3. Generalization Ability of the Learned Policy
We randomize the initial positions of the aircrafts to evaluate the generalization ability of the proposed learning framework. To investigate the generalization ability of the learned policy, we introduce random offsets in the evaluation phase. The random perturbations range from 00 to $100$ meters in the x, y, and z dimensions of the initial Cartesian coordinates of the aircraft to simulate uncertainty factors in real-world conditions. According to the experimental results in Table 4, our method has demonstrated outstanding performance, with the adaptive weighting method achieving a remarkable hit success rate of $98\%$. By contrast, although the linear weighting method also performed well, its generalization performance was not as good as the E-SAC method.
Table 4. Experiment results in the random initialization setting
Algorithm Success Rate Hit success rate Return
Ours (adaptive) $\bf{98.0\%\pm 1.3\%}$ $\bf{98.0\%\pm 1.3\%}$ $\bf{-1436.0\pm 238.9}$
Ours (linear) $86.0\%\pm 5.4\%$ $86.0\%\pm 5.4\%$ $-5800.8\pm 1420.3$
TD3 $0.0\%\pm 0.0\%$ $0.0\%\pm 0.0\%$ $-5720.9\pm 715.8$
E-SAC $90.0\%\pm 2.8\%$ $90.0\%\pm 2.8\%$ $-3722.2\pm 395.5$
SAC $44.0\%\pm 3.3\%$ $0.0\%\pm 0.0\%$ $-8318.1\pm 822.8$
BC $22.4\%\pm 3.2\%$ $22.4\%\pm 3.2\%$ $-20504.7\pm 1156.3$
To test the generalization ability and the launch efficiency, we further evaluate the models in the unlimited missile supply setting. Note that in the training phase, the aircraft only carry one missile. Table 5 shows the launch efficiency (calculated by the fraction between the hit number and the launch number) of the four methods that have mastered the missile launch policy in the unlimited missile setting, and the methods that cannot effectively launch are omitted. The proposed learning framework has exhibited outstanding launch efficiency in this experiment, accurately launching missiles regardless of the lock conditions, which demonstrates the effectiveness and reliability of the learned launch policy.
Table 5. Launch efficiency in the unlimited missile setting
@@ 6.3. Generalization Ability of the Learned Policy
Ours (adaptive) $\bf{100\%}$
Ours (linear) $\bf{100\%}$
E-SAC $11.4\%$
BC $92.3\%$
6.4. Ablation Study on Expert Data
To evaluate the impact of expert data quality on the learning performance of our framework, we conduct ablation studies using two types of expert data: one collected in AI mode and the other collected by the BC policy. By using the expert data collected by the AI mode, the proposed framework can achieve an average return of $-483.68$, which is a little higher than that of employing the BC data as the expert demonstration. The results are shown in Figure 9. The experiment results demonstrate that the performance of using AI data and BC data only have a slight difference. This phenomenon indicates that the proposed framework is robust to the quality of expert data. Although the lower quality of BC expert data results in instability during the training process and fluctuations in success rates, the highest success rates of both modes have reached $100\%$, which validates the effectiveness of the proposed framework.
Figure 9. Ablation studies on expert data in the proposed framework.
@@ 7. Conclusion
This study focuses on the complex multi-dimensional decision-making problems in UCAV dogfight. Leveraging the Harfang3D sandbox, we construct a highly realistic simulation platform and propose an innovative high-imitative reinforcement learning framework for dogfight policy. This method integrates expert experience with autonomous exploration mechanisms, achieving efficient and precise policy learning. Experimental results demonstrate that the constructed simulation environment serves as an effective platform for conducting reinforcement learning research in aerial combat, enabling UCAVs to handle complex, multistage dogfight tasks, including target pursuit, target lock, and missile launch in simulated air combat missions. Policies learned with the proposed framework can achieve a hit success rate of up to $100.0\%$. Compared to RL and imitation learning algorithms, the learned policies in this study demonstrate superior success rates and robustness while requiring less expert data.
Although our framework effectively performs multistage tasks, it exhibits a certain dependence on the quality of expert data, and its training outcomes are influenced by the reward design. In the future, we aim to explore a hierarchical policy to gather representative knowledge from limited expert data, striving to reduce the reliance on expert data quality.

### v2

@@ 6.1. Comparative Results
According to Table 3, the adaptive weighting method proposed in this study demonstrates superior performance across two key measures (success rates and returns) against three opponent modes. Specifically, the TD3 and DSAC-v2 methods show the lowest performance in terms of the hit success rate, which is essentially $0\%$, among all methods. In contrast, the method proposed in this study, along with the SAC method, achieve a relatively high hit success rate. Further comparison between our method and the BC method reveals that, the BC method performs more poorly in terms of both success rate in task completion and reward value compared to our method due to the lack of interaction with the environment and the inability to explore autonomously. It is noteworthy that, although the BC method achieves a higher success rate than the TD3 and DSAC-v2 methods in the serpentine maneuvering and circling maneuvering opponent modes, its average returns are significantly lower. This is also attributed to the inability to explore autonomously. When compounding errors occur (xu2020error, ; ross2010efficient, ; ross2011reduction, ), BC strategy tends to lose control and thus achieves very low returns, whereas TD3 and DSAC-v2 methods can still closely track the opponent by pursuit strategy, thus preventing a rapid decline in returns. Comparing the first row and the last row, the proposed approach, benefiting from autonomous exploration, has even accomplished better performance than the expert dataset.
@@ 6.2. Analytical Results
We show weight $\lambda$ in the two modes (linear and adaptive) in Figure 6, and find that the value of $\lambda$ is related to the learning performance. As shown in Figure 5, the linear weighting method proposed in this study demonstrates a faster learning rate in the initial stages of training compared to that of the adaptive weighting method. This is primarily attributed to the higher weight assigned to the imitation of objectives by the linear weighting method in the early stages, accelerating the learning of expert knowledge. However, this approach increases the risk of overfitting and limits the model’s generalization ability, resulting in instability and poor convergence performance in training processes. Since we show the average $\lambda$ over a batch data in Figure 6, the $\lambda$ in the adaptive mode is decimal between 0 and 1.
@@
Figure 7. The distance between our aircraft and the opponent in the serpentine opponent mode.
@@
Figure 7 depicts the distance between our aircraft and the opponent in one episode, and the trajectories are collected by the convergent models learned by the corresponding algorithms in the serpentine opponent mode. Subfigures in Figure 7 annotate the moments of missile launch (red triangles), phases of holding missiles (blue background), and stages of successfully locking onto targets (yellow background). If the first red dot on the curve falls within the intersection of the blue and yellow backgrounds, it indicates that the agent has mastered an effective launch policy. Accordingly, it is observed that the method proposed, along with the SAC, E-SAC, and BC methods, successfully executed missile launches and accurately shot down the opponent aircraft (i.e., mastering the three-dimensional policy of pursuit, lock, and launch). Further comparison of the speed of target lock and missile launch among these four methods reveals that the proposed method achieves both with fewer steps, indicating higher attack efficiency. The distance curves of the TD3 and DSAC-v2 methods exhibit a downward trend, and there are target locking zones; however, they fail to correctly launch missiles, demonstrating their mastery of only the two-dimensional policy of pursuit and lock. Although there is a probability of successful firing in the optimal results of TD3, it does not converge and lacks stability; therefore, we do not consider it to have learned the launch policy. Furthermore, although neither TD3 nor DSAC-v2 can correctly launch missiles, they occasionally manage to succeed by colliding with the opponent during testing, indicating that their pursuit and locking strategies are relatively well-learned.
Figure 8. The trajectories collected by the learned policy (blue) and the opponent (red) in the serpentine opponent mode. The solid lines represent the 3D trajectories, while the dashed lines are their projections onto the horizontal plane. The lighter color denotes a smaller timestep, and the darker color denotes approaching the end of the episode. The red triangles on the blue trajectories indicate the moments of launch.
@@
Figure 8 corresponds respectively to Figure 7, and Figure 8 shows the three-dimensional interaction scenarios between our aircraft and the opponent aircraft using three different methods. The red curves represent the movement trajectories of the opponent aircraft, while the blue curves represent the movement trajectories of our aircraft, with the colors gradually deepening over time. The moments of launch are indicated by red triangles on the blue trajectories, and the yellow areas on the blue trajectories represent the stages of successfully locking onto the opponent. Figure 8(a) illustrates the operational effectiveness of the proposed learning framework, i.e., the controlled UCAV swiftly catches up to the opponent aircraft, successfully locking onto it, launching a missile, and accurately hitting the target. Note that the blue curve is nearly the shortest path from the start position of our aircraft to the opponent. In contrast, the UCAV with the TD3 (Figure 8(c)) or DSAC-v2 (Figure 8(f)) policy can pursue the opponent aircraft but fails to shoot it down due to premature missile launches, indicating insufficient launch capabilities. Through the above analysis, we systematically summarize the policies learned by all the methods, presented in Table 4.
Figure 9. The trajectories collected by the policy learned by the proposed method (blue) and the opponent (red) in three opponent modes. The solid lines represent the 3D trajectories, while the dashed lines are their projections onto the horizontal plane. The lighter color denotes a smaller timestep, and the darker color denotes approaching the end of the episode. The red triangles on the blue trajectories indicate the moments of launch.
In addition to the serpentine mode, we also plotted the results of our method in other opponent modes, as shown in Figure 9. These subfigures demonstrate that the proposed approach can work against various opponent types, shooting down the opponent aircraft in a fast way.
Table 4. Policy analysis.
@@
6.3. Robustness of the Learned Policy
To evaluate the robustness and launch efficiency, we evaluate the models in the unlimited missile supply setting in the serpentine opponent mode. Note that in the training phase, the aircraft only carry one missile. Table 5 shows the launch efficiency (calculated by the fraction between the hit number and the launch number) of the five methods that have mastered the missile launch policy in the unlimited missile setting, and the methods that cannot effectively launch are omitted. The proposed learning framework has exhibited outstanding launch efficiency in this experiment, accurately launching missiles regardless of the lock conditions, which demonstrates the effectiveness and reliability of the learned launch policy.
Table 5. Launch efficiency in the unlimited missile setting.
@@ 6.3. Robustness of the Learned Policy
Ours (adaptive) $\bf{98.8\%}$
Ours (linear) $98.5\%$
E-SAC $71.1\%$
SAC $69.0\%$
BC $94.9\%$
@@ 7. Conclusion
This study focuses on the complex multi-dimensional decision-making problems in the UCAV dogfight. Leveraging the Harfang3D sandbox, we construct a realistic simulation platform and propose an innovative high-imitative reinforcement learning framework for dogfight policy. This method integrates expert experience with autonomous exploration mechanisms, achieving efficient and precise policy learning. Experimental results demonstrate that the constructed simulation environment serves as an effective platform for conducting reinforcement learning research in aerial combat, enabling UCAVs to handle complex, multistage dogfight tasks, including target pursuit, target lock, and missile launch in simulated air combat missions. Policies learned with the proposed framework can achieve a hit success rate of up to $100.0\%$. Compared to RL and imitation learning algorithms, the learned policies in this study demonstrate superior success rates and robustness while requiring less expert data.
Our framework can effectively complete multistage tasks, but it exhibits a certain dependence on the quality of expert data, and its training outcomes are influenced by the reward design. In the future, we aim to explore a hierarchical policy to gather representative knowledge from limited expert data, striving to reduce the reliance on expert data quality.